# Magneto-optical transport in tilted type-I double-Weyl semimetals in the presence of orbital magnetic moment

### v1

@@ III.2 Calculations of conductivities components linear in magnetic field
Substituting Eq.(III) with B terms up to first order into the first term of Eq.(III), we get conductivities components
@@
$\sigma_{ab}^{(B,\Omega)}(\omega)=\frac{\tau}{(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k[(\mathit{v}_{a}^{s}B_{b}+\mathit{v}_{b}^{s}B_{a})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})-\mathit{v}_{a}^{s}\mathit{v}_{b}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (23)
$\sigma_{ab}^{(B,m)}(\omega)=\frac{\tau}{(1-i\omega\tau)}\frac{e^{2}}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\frac{\partial\mathit{v}_{a}^{s}}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}\mathit{v}_{b}^{s}\Bigr{]}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (24)
We can easily check from Eqs.(22) and (23) that $\sigma_{ab}^{(B)}(\omega)=\sigma_{ba}^{(B)}(\omega)$. The system possesses the time-reversal symmetry without tilt term and therefore condutivities will vanish [22, 28]. However, for finite value of tilt $t_{s}\neq 0$, the time-reversal symmetry is broken. The first term in Eq. (22) is related to the Berry curvature $\Omega$. The second term in Eq. (22) stems from the contribution of the orbital magnetic moment for isotropic WSMs [22]. In the following, the detailed analysis of Eq. (22) is given by considering the magnetic field B perpendicular and parallel to the tilt directon $t_{s}$.
@@
In this case, Hall conductivities components are zero and one can get following expressions for longtudinal components
$\sigma_{zz}^{(B,\Omega)}=\frac{\hbar\pi\mathit{v}_{F}^{2}\sigma_{1}^{(B)}s}{3\pi\alpha_{2}\mu}\Bigl{[}\frac{3}{2}\frac{1-2t_{s}^{2}-t_{s}^{4}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}$
$-\frac{(-3+5t_{s}^{2}-3t_{s}^{3}+3t_{s}^{4})}{t_{s}^{3}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xx}^{(B,\Omega)}=\sigma_{1}^{(B)}s\Bigl{[}\frac{(-4+4\sqrt{1-t_{s}^{2}}+t_{s}^{2}(3-\sqrt{1-t_{s}^{2}})}{t_{s}^{5}}\Bigr{]}$
where $\sigma_{1}^{(B)}=\frac{e^{2}\tau\mu eB\alpha_{2}}{\hbar(1-i\omega t)2\pi\mathit{v}_{F}\hbar^{2}}$.
$\sigma_{zz}^{(B,m)}(\omega)$ $=$ $\frac{\hbar\mathit{v}_{F}^{2}\sigma_{0}^{(B)}s}{\pi\alpha_{2}\mu}\Bigl{[}\frac{-3+5t_{s}^{2}}{3t_{s}^{3}}-\frac{(1-t_{s}^{2})^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xx}^{(B,m)}(\omega)$ $=$ $\sigma_{0}^{(B)}s\Bigl{[}\frac{(-4-2t_{s}^{4}+4\sqrt{1-t_{s}^{2}}+t_{s}^{2}(7-5\sqrt{1-t_{s}^{2}})}{t_{s}^{5}\sqrt{1-t_{s}^{2}}}\Bigr{]}$ (28)
where $\sigma_{1}^{(B)}=\frac{e^{2}\tau\mu eB\alpha_{2}}{\hbar(1-i\omega t)2\pi\mathit{v}_{F}\hbar^{2}}$.
From the Eqs.(III.2) and (III.2), we have obtained non-trivial dependence of conductivities components on tilt parameter.
@@
We can represent the magnetic field in x-y plane as $\textbf{B}=B(\hat{i}\cos\gamma+\hat{j}\sin\gamma)$ where $\gamma$ is the angle between magnetic field B and x-axis. In this case longitudinal components of condutivities are zero and we get the following expressions for planer Hall conductivities [22, 29]
@@
$\sigma_{yz}^{(B)}(\omega)=[2\sigma^{(B,\Omega)}(\omega)+\sigma^{(B,m)}(\omega)]\sin\gamma$ (30)
@@
$\sigma^{(B,\Omega)}_{xz}(\omega)=\frac{\tau}{(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k[\mathit{v}_{z}^{s}B\cos\gamma(\mathit{v}_{x}^{s}\Omega_{ky}+\mathit{v}_{z}^{s}\Omega_{kz})-\mathit{v}_{z}^{s}B\sin\gamma\mathit{v}_{x}^{s}\Omega_{ky}]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (31)
@@
$\sigma^{(B,m)}_{xz}(\omega)=\frac{\tau}{(1-i\omega\tau)}\frac{e^{2}}{(2\pi)^{3}}\int d^{3}k[B\cos\gamma(\frac{\partial\mathit{v}_{x}^{s}}{\partial k_{z}}m_{kx}^{s}-v_{z}^{s}\frac{\partial\mathit{m}_{kx}^{s}}{\partial k_{z}})-B\sin\gamma(\frac{\partial\mathit{v}_{x}^{s}}{\partial k_{z}}m_{ky}^{s}-v_{z}^{s}\frac{\partial\mathit{m}_{ky}^{s}}{\partial k_{x}})]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (32)
The second term of both the above equations contribute to zero.
@@
$\sigma^{(B,\Omega)}=\sigma_{0}^{(B)}s\Bigl{[}\frac{-3t_{s}+5t_{s}^{3}-6t_{s}^{5}}{t_{s}^{4}}-\frac{3}{2}\frac{(1-t_{s}^{2})^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma^{(B,m)}=\sigma_{0}^{(B)}s\Bigl{[}\frac{(3-2t_{s}^{2})}{t_{s}^{3}}+\frac{3}{2}\frac{(1-t_{s}^{2})^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
@@
where $\sigma_{0}^{(B)}=\frac{e^{3}\tau B\mathit{v}_{F}}{\hbar(1-i\omega t)12\pi^{2}\hbar}$. From Eqs. (24)–(33), we notice that the B-linear magnetoconductivity in tilted Weyl semimetals is independent of Fermi energy $\mu$, and the odd function of $t_{s}$ as in the case of single WSMs [22]. According to the Nielsen-Ninomiya theorem [5, 5], the Weyl nodes with opposite chirality always appear in pairs. The total magnetoconductivity of the system is the sum of all the Weyl nodes. Therefore, for the case of $t_{+}=t_{-}$ , where the tilt inversion symmetry is broken, the contribution of the Weyl node to the magnetoconductivity has the opposite sign for the opposite (B) chirality , giving rise to $\sigma_{ab}(\omega)$ = 0. Whereas, for the case of $t_{+}=-t_{-}$ , where the tilt inversion symmetry is unbroken,each Weyl node produces an identical contribution to the magnetoconductivity, and so the nonzero magnetoconductivity emerges for this case.
III.3 Calculatons of quadractic-B contributon to the conductivity $\sigma_{ab}^{(B^{2})}$

### v2

@@ III.2 Calculations of conductivities components linear in magnetic field
Substituting Eq.(16) with B terms up to first order into the first term of Eq.(17), we get conductivities components
@@
$\sigma_{ab}^{(B,\Omega)}(\omega)$ $=$ $\frac{\tau}{\hbar(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k[(\mathit{v}_{a}^{s}B_{b}+\mathit{v}_{b}^{s}B_{a})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})-\mathit{v}_{a}^{s}\mathit{v}_{b}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (27)
$\sigma_{ab}^{(B,m)}(\omega)$ $=$ $\frac{\tau}{\hbar(1-i\omega\tau)}\frac{e^{2}}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\frac{\partial\mathit{v}_{a}^{s}}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}\mathit{v}_{b}^{s}\Bigr{]}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (28)
We can easily check from Eqs.(27) and (28) that $\sigma_{ab}^{(B)}(\omega)=\sigma_{ba}^{(B)}(\omega)$. The system possesses the time-reversal symmetry without tilt term and therefore conductivities will vanish [22, 32]. However, the time-reversal symmetry is broken for a finite value of tilt $t_{s}\neq 0$. The first term and the second term in Eq. (27) is related to the Berry curvature $\Omega$ and the orbital magnetic moment respectively and has been studued in details for isotropic WSMs [22]. We will explore our study for double and triple-WSMs case. In the following, the detailed analysis of Eq. (27) is given by considering the magnetic field B perpendicular and parallel to the tilt direction $t_{s}$.
Refer to caption  Refer to caption  Refer to caption
Figure 4: The frequency dependence of optical conductivity at t=0.5 and B = 1 T for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
@@
In this case, Hall conductivities components are zero, and one can get the following expressions for longitudinal components.
for J=1
$\sigma_{zz}^{(B,\Omega)}$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{2(3-5t_{s}^{2}-3t_{s}^{4})}{3t_{s}^{2}}-\frac{(t_{s}^{2}-1)^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xx}^{(B,\Omega)}$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{2t_{s}^{2}-3}{3t_{s}^{3}}-\frac{1-t_{s}^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}B\Bigr{]}$ (30)
for J=2
$\sigma_{zz}^{(B,\Omega)}$ $=$ $\frac{2\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{2(3-5t_{s}^{2}-3t_{s}^{4})}{3t_{s}^{2}}-\frac{(t_{s}^{2}-1)^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xx}^{(B,\Omega)}=\frac{\mu\alpha_{2}\sigma_{1}^{(B)}s}{\pi\mathit{v_{F}\hbar^{2}}}\Bigl{[}\frac{(-4+4\sqrt{1-t_{s}^{2}}+t_{s}^{2}(3-\sqrt{1-t_{s}^{2}})}{t_{s}^{5}}\Bigr{]}$
for J= 3
$\sigma_{zz}^{(B,\Omega)}$ $=$ $\frac{3\mathit{v}_{F}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{2(-3+5t_{s}^{2}+3t_{s}^{4})}{3t_{s}^{3}}-\frac{(-1+t_{s}^{2})^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (33)
$\sigma_{xx}^{(B,\Omega)}$ $=$ $\frac{\mu^{\frac{4}{3}}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B)}s}{4\pi^{\frac{3}{2}}\mathit{v}_{F}\hbar^{\frac{7}{3}}}\Bigl{[}-\frac{(1152(3)^{\frac{1}{2}}\Gamma(\frac{17}{6})({{}_{2}F_{1}}(\frac{7}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})-{{}_{2}F_{1}}(\frac{7}{3},\frac{11}{3},\frac{19}{3},\frac{2t_{s}}{-1+t}))}{143(1-t_{s})^{\frac{7}{3}}\Gamma(\frac{10}{3})}\Bigr{]}$ (34)
where $\sigma_{1}^{(B)}=\frac{e^{3}\tau B}{(1-i\omega\tau)2\hbar}$.
So, we can rewrite the above components follow the power laws $\sigma_{zz}^{(B,\Omega)}(\omega)\propto J$ and $\sigma_{xx}^{(B,\Omega)}(\omega)\propto\frac{\mu^{2\bigl{(}1-\frac{1}{J}\bigr{)}}\alpha_{J}^{2/J}}{\hbar^{\bigl{(}3-\frac{2}{J}\bigr{)}}}$.
Via the similar calculation, Eq.(28) becomes for J=1
$\sigma_{zz}^{(B,m)}(\omega)$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{2(-3+5t_{s}^{2})}{3t_{s}^{3}}-\frac{(t_{s}^{2}-1)^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xx}^{(B,m)}(\omega)$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{(3-8t_{s}^{2})}{3t_{s}^{3}}-\frac{1-3t_{s}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (36)
for J=2
$\sigma_{zz}^{(B,m)}(\omega)$ $=$ $\frac{\sigma_{1}^{(B)}s}{3\pi^{2}}\Bigl{[}\frac{-3+5t_{s}^{2}}{3t_{s}^{3}}-\frac{(1-t_{s}^{2})^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (37)
$\sigma_{xx}^{(B,m)}(\omega)$ $=$ $\frac{\mu\alpha_{2}\sigma_{1}^{(B)}s}{\pi\mathit{v}_{F}\hbar^{2}}\Bigl{[}\frac{(-4-2t_{s}^{4}+4\sqrt{1-t_{s}^{2}}+t_{s}^{2}(7-5\sqrt{1-t_{s}^{2}})}{t_{s}^{5}\sqrt{1-t_{s}^{2}}}\Bigr{]}$
for J=3
$\sigma_{zz}^{(B,m)}(\omega)$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{2\pi^{2}\hbar}\Bigl{[}\frac{(-3+5t_{s}^{2})}{t_{s}^{3}}-\frac{3(-1+t_{s}^{2})^{2}}{2}\ln{\frac{1-t_{s}}{1+t_{s}}}\Bigr{]}$ (39)
$\sigma_{xx}^{(B,m)}(\omega)$ $=$ $\frac{7\alpha_{3}^{\frac{2}{3}}\mu^{\frac{4}{3}}\sigma_{1}^{(B)}s}{\pi^{\frac{1}{2}}\mathit{v}_{F}\hbar^{\frac{4}{3}}}\Bigl{[}\frac{\frac{5(-65-39t_{s}+93t_{s}^{2})(1-t_{s})^{\frac{8}{3}}}{t_{s}^{3}(1+t_{s})^{\frac{2}{3}}}+243\frac{(1-t_{s})^{5/3}}{(1+t_{s})^{2/3}}-\frac{5(1+t_{s})(-65+126t_{s}^{2}){}_{2}F_{1}(\frac{7}{3},\frac{8}{3},\frac{10}{3},\frac{2t_{s}}{-1+t_{s}})}{t_{s}^{3}}}{243(3)^{\frac{1}{2}}(1-t_{s})^{\frac{7}{3}}(1+t_{s})\Gamma(\frac{13}{6})(\frac{7}{3})}\Bigr{]}$
where $\sigma_{1}^{(B)}=\frac{e^{3}\tau B}{(1-i\omega\tau)2\hbar}$.
So, we can summarize the above components follow the power laws $\sigma_{zz}^{(B,m)}(\omega)\propto J$ and $\sigma_{xx}^{(B,m)}(\omega)\propto\frac{\mu^{2\bigl{(}1-\frac{1}{J}\bigr{)}}\alpha_{J}^{2/J}}{\hbar^{2(\bigl{(}1-\frac{1}{J}\bigr{)}}}$.
@@
We can represent the magnetic field in the x-y plane as $\textbf{B}=B(\hat{i}\cos\gamma+\hat{j}\sin\gamma)$ where $\gamma$ is the angle between magnetic field B and x-axis. In this case, longitudinal components of conductivities are zero and we get the following expressions for planer Hall conductivities [22, 33]
@@
$\sigma_{yz}^{(B)}(\omega)=[\sigma^{(B,\Omega)}(\omega)+\sigma^{(B,m)}(\omega)]\sin\gamma$ (42)
@@
$\sigma^{(B,\Omega)}_{xz}(\omega)=\frac{\tau}{\hbar(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k[\mathit{v}_{z}^{s}B\cos\gamma(\mathit{v}_{x}^{s}\Omega_{ky}+\mathit{v}_{z}^{s}\Omega_{kz})-\mathit{v}_{z}^{s}B\sin\gamma\mathit{v}_{x}^{s}\Omega_{ky}]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (43)
@@
$\sigma^{(B,m)}_{xz}(\omega)=\frac{\tau}{\hbar(1-i\omega\tau)}\frac{e^{2}}{(2\pi)^{3}}\int d^{3}k\Bigl{[}B\cos\gamma\Bigl{(}\frac{\partial\mathit{v}_{x}^{s}}{\partial k_{z}}m_{kx}^{s}-v_{z}^{s}\frac{\partial\mathit{m}_{kx}^{s}}{\partial k_{z}}\Bigr{)}-B\sin\gamma\Bigl{(}\frac{\partial\mathit{v}_{x}^{s}}{\partial k_{z}}m_{ky}^{s}-v_{z}^{s}\frac{\partial\mathit{m}_{ky}^{s}}{\partial k_{x}}\Bigr{)}\Bigr{]}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (44)
The second term of both the above equations contributes to zero.
for J=1
$\sigma^{(B,\Omega)}$ $=$ $\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{-3+5t_{s}^{2}-6t_{s}^{4}}{3t_{s}^{3}}-\frac{(1-t_{s}^{2})^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma^{(B,m)}=\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}-\frac{2t_{s}^{2}-3}{3t_{s}^{3}}+\frac{1-t_{s}^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (46)
for J=2
$\sigma^{(B,\Omega)}$ $=$ $2\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{-3t_{s}+5t_{s}^{3}-6t_{s}^{5}}{3t_{s}^{4}}-\frac{(1-t_{s}^{2})^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma^{(B,m)}=3\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{6\pi^{2}\hbar}\Bigl{[}\frac{(3-2t_{s}^{2})}{3t_{s}^{3}}+\frac{(1-t_{s}^{2})}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (48)
for J=3
@@
$\begin{split}\sigma^{(B,\Omega)}=&3\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{(-3t_{s}+5t_{s}^{3}-6t_{s}^{5})}{3t_{s}^{4}}\\ &-\frac{(-1+t_{s}^{2})^{2}}{2t_{s}^{4}}\ln{\frac{1-t_{s}}{1+t_{s}}}\Bigr{]}\end{split}$ (49)
$\sigma^{(B,m)}=3\frac{\mathit{v_{F}}\sigma_{1}^{(B)}s}{4\pi^{2}\hbar}\Bigl{[}\frac{(3-2t_{s}^{2})}{3t_{s}^{3}}-\frac{(-1+t_{s}^{2})}{2t_{s}^{4}}\ln{\frac{1-t_{s}}{1+t_{s}}}\Bigr{]}$ (50)
where $\sigma_{0}^{(B)}=\frac{e^{3}\tau B}{(1-i\omega\tau)2\hbar}$. So, we can rewrite the above components follow the power laws $\sigma^{(B,\Omega)}\propto\frac{J}{\hbar}$ and $\sigma^{(B,m)}\propto\frac{J}{\hbar}$.
@@
From Eqs. (33)–(50), we notice that the B-linear magnetoconductivity in tilted mult-Weyl semimetals is independent of Fermi energy $\mu$, and the odd function of $t_{s}$ as in the case of single WSMs [22]. According to the Nielsen-Ninomiya theorem [5, 5], the Weyl nodes with opposite chirality always appears in pairs. The total magnetoconductivity of the system is the sum of all the Weyl nodes. Therefore, for the case of $t_{+}=t_{-}$, where the tilt inversion symmetry is broken, the contribution of the Weyl node to the magnetoconductivity has the opposite sign for the opposite (B) chirality, giving rise to $\sigma_{ab}(\omega)$ = 0. Whereas, for the case of $t_{+}=-t_{-}$, where the tilt inversion symmetry is unbroken, each Weyl node produces an identical contribution to the magnetoconductivity, and so the nonzero magnetoconductivity emerges for this case. Figure (3) shows the B-linear magnetoconductivity as a function of the tilt $t_{+}$ . It is observed that the contribution from the orbital magnetic moment is always negative and decreases with increasing $t_{+}$, which partially cancels the contribution of the Berry curvature to magnetoconductivity [22]. The total magnetoconductivity tends to slow down at the large $t_{+}$ . Figure (4) shows the B-linear magnetoconductivity as a function of the THz incident light.
III.3 Calculations of quadratic-B contribution to the conductivity $\sigma_{ab}^{(B^{2})}$